OpenStudy (anonymous):

solve. 3(n-2) - 2(3-n)= 4(n-3)

OpenStudy (anonymous):

so you need to use the distributive property to get: \[3n - 6 - 6 + 2n = 4n - 12\]

OpenStudy (anonymous):

you can combine terms to get: \[5n - 12 = 4n - 12\]

OpenStudy (anonymous):

you can add 12 to both sides to get: \[5n = 4n\]

OpenStudy (anonymous):

you can subtract 4n from both sides to get: \[n = 0\]

OpenStudy (anonymous):

if you plug it back in: \[3(-2) -2(3) = -6 - 6 = -12 = 4(-3)\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome